OpenStudy (anonymous):

Find three solutions of the equation. y=3x+4 a. (–2, –2), (1, 7), (2, 11) c. (–2, –2), (1, 7), (0, 4) b. (0, 4), (3, 11), (–2, –2) d. (1, 7), (0, 4), (–1, 0)

OpenStudy (anonymous):

plug in the 1st number of the ordered pair in for x and solve, if it solves to the 2nd number in the ordered pair then ut is a solution

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@charlotte123

OpenStudy (luigi0210):

Char's got dis.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks u very much

OpenStudy (charlotte123):

Thanks to my handy dandy Calculator - It turns out that the Second Option Is correct :) -2 = 3(-2) + 4 <---- As 3(-2) = -6 and -6 + 4 = -2! 7 = 3(1) + 4 <--- As 3(1) = 3 and 3 + 4 = 7! 4 = 3(0) + 4 <---- As 3(0) = 0 and 0 + 4 = 4! Using the Equation ----> y = 3x + 4 <----

OpenStudy (charlotte123):

@kmacdaddy No problem! Glad I helped!!!

OpenStudy (charlotte123):

@Luigi0210 (⌐■_■)

OpenStudy (anonymous):

@Luigi0210 I cant send u a message why is that?

OpenStudy (anonymous):

@luigi hates every1. thats y